OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (-2, 0) and (2, 8)?

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

y2-y1/x2-x1=Slope so 8-0/2-(-2)

OpenStudy (anonymous):

you can find the slope of the line by following formula. \[m= \frac{ y_{2}- y_{1} }{ x_{2}- x_{1} } = \frac{ 8-0 }{ 2-(-2) } = \frac{ 8 }{ 4 } = 2\] and now equation is \[(y-y _{1}) = m (x-x _{1}) \] \[(y-0) = 2 (x + 2) \implies y = 2x+ 4 \]

OpenStudy (anonymous):

same copy @Smg93

OpenStudy (anonymous):

so would it be y=-2(x-2)?

OpenStudy (anonymous):

no @Tfinley2013 you can crosscheck by substituting points in \[y = 2x + 4\] it will satisfy, but it wont satisfy your equation.

OpenStudy (anonymous):

so would it be y=2(x+2) ?